A sample of krypton gas at a pressure of 797 mm Hg and a temperature of 87°C, occupies a volume of 8.10 liters. If the gas is heated at constant pressure to a temperature of 125 °C, the volume of the gas sample will be ? L.
Answer #1


here at constant pressure means no change in pressure but change in temperature and volume

by using the charle's law formula V1 / T1 = V2 / T2 and then calculate the mew volume

here the values are => V1 = 8.10 L     ,    T1 = 87 c = 360.15 K

V2 = ? , T2 = 125 c = 398.15 K

substituting the values in the equation

V2 = V1 T2 / T1

V2 = 8.10 * 398.15 / 360.15

V2 = 8.9546 liters

answer => 8.95 L